# Ultra-sensitive sizing of individual nanoparticles with an optofluidic microcavity

### v1

@@ II Conclusions
In this work, we have demonstrated a technique for measuring and analyzing single, unlabeled, freely diffusing nanoparticles down to a diameter of 3 nm, which corresponds to about 20 gold atoms across and an effective molecular weight of 160 kDa. We note that by operating far away from the plasmon resonance wavelength of gold nanospheres ($\approx$ 500 nm), we can neglect plasmon enhancement, and the GNPs behave essentially like dielectric particles. Our smallest GNPs have $\Re(\alpha)$ equivalent to a silica sphere with a diameter of about 8 nm. This approaches the single-molecule sensitivity recently reported with a photothermally enhanced optical microcavity [17], while additionally offering the capability for precise size determination. Using the modified autocorrelation model described in this work, it is possible to quantitatively extract information about single particles from the autocorrelation of their event traces without the need for a priori calibration measurements.
@@
Furthermore, the high measurement bandwidth possible with this method provides a temporal resolution orders of magnitude better than imaging-based nanoscopy techniques [5], and can enable not only the resolution of the motion of smaller and thus faster particles, but also of fast dynamics such as rotation or internal motion such as conformational dynamics in the future.
III Methods
@@
Experimental setup. The cavity fibers were machined in-house with single-mode fibers according to the procedure developed by Hunger et al. [19] and the profiles had radii of curvature of 43 µm and 41 µm. The mirror coating was done by Laseroptik GmbH. The cavity is probed with a grating-stabilized diode laser at 780 nm. Transmitted light detected by an avalanche photodetector is monitored on an oscilloscope and used as the error signal for cavity locking.
@@
For cavity locking, we implement a slow stabilization scheme designed such that rapid resonance shifts due to interaction with a diffusing nanoparticle lie outside the locking bandwidth, pass out of the cavity unattenuated and are detected in the transmission signal. Slower acoustic and mechanical jitter as well as thermal drifts are, on the other hand, corrected by the locking system. A detailed verification of this is shown in Sec. S1.
@@
Data processing. Data measured by the oscilloscope were smoothed with a Savinsky-Golay filter and resampled to 500 kHz (5 nm - 20 nm nanospheres) or 1 MHz (3 nm nanospheres). Frequency filtering was performed to remove sharp peaks arising from noise sources. The data were then processed as described in Sec. I.4 by a custom-written Python script. The autocorrelation was computed using the multiple-tau algorithm [24]. For details about the data processing steps, see Supplementary Information.
III.1 Acknowledgments
We acknowledge Lucas Flatten for helpful discussions and Jens Treptow for performing TEM measurements. DH and SP acknowledge funding from the Max Planck School of Photonics (MPSP) and the Karlsruhe School of Optics and Photonics (KSOP). DH, SP, CR and CF acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Collaborative Research Centre “4f for Future” (CRC 1573 project number 471424360, project C4).
III.2 Competing interests
@@ III.3 Author contributions
III.4 Additional information
Supplementary information is available for this paper.
Correspondence and requests for materials should be addressed to David Hunger or Shalom Palkhivala.
Reprints and permissions information is available at www.nature.com/reprints.

### v2

@@ 3 Conclusions
In this work, we have demonstrated a technique for measuring and analyzing single, unlabeled, freely diffusing nanoparticles down to a diameter of 3 nm, which corresponds to about 20 gold atoms across and an effective molecular weight of 160 kDa. We note that by operating far away from the plasmon resonance wavelength of gold nanospheres ($\approx$ 500 nm), we can neglect plasmon enhancement, and the GNPs behave essentially like dielectric particles. Our smallest GNPs have $\Re(\alpha)$ equivalent to a silica sphere with a diameter of about 8 nm. This approaches the single-molecule sensitivity recently reported with a photothermally enhanced optical microcavity 17, while additionally offering the capability for precise and accurate size determination. Using the modified autocorrelation model described in this work, it is possible to quantitatively extract information about single particles from the autocorrelation of their event traces without the need for a priori calibration measurements.
@@
Furthermore, the high measurement bandwidth possible with this method provides a temporal resolution orders of magnitude better than imaging-based nanoscopy techniques 5, and can enable not only the resolution of the motion of smaller and thus faster particles, but also of fast dynamics such as rotation or internal motion such as conformational dynamics in the future.
4 Methods
@@
Experimental setup. The cavity fibers were machined in-house with single-mode fibers according to the procedure developed by Hunger et al. 19 and the profiles had radii of curvature of 43 µm and 41 µm. The mirror coating was done by Laseroptik GmbH. The cavity is probed with a grating-stabilized diode laser at 780 nm. Transmitted light detected by an avalanche photodetector is monitored on an oscilloscope and used as the error signal for cavity locking.
@@
For cavity locking, we implement a slow stabilization scheme designed such that rapid resonance shifts due to interaction with a diffusing nanoparticle lie outside the locking bandwidth, pass out of the cavity unattenuated and are detected in the transmission signal. Slower acoustic and mechanical jitter as well as thermal drifts are, on the other hand, corrected by the locking system. A detailed verification of this is shown in the Supporting Information, see Figs. S3 and S4.
@@
Data processing. Data measured by the oscilloscope were smoothed with a Savinsky-Golay filter and resampled to 500 kHz (5 nm - 20 nm nanospheres) or 1 MHz (3 nm nanospheres). Frequency filtering was performed to remove sharp peaks arising from noise sources. The data were then processed as described above by a custom-written Python script. The autocorrelation was computed using the multiple-tau algorithm 24. For details about the data processing steps, see Supporting Information.
4.1 Acknowledgments
We acknowledge L. Flatten for helpful discussions and J. Treptow for performing TEM measurements. D.H. and S.P. acknowledge funding from the Max Planck School of Photonics (MPSP) and the Karlsruhe School of Optics and Photonics (KSOP). D.H., S.P., C.R. and C.F. acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Collaborative Research Centre “4f for Future” (CRC 1573 project number 471424360, project C4).
4.2 Competing interests
@@ 4.3 Author contributions
4.4 Supporting information
The following supporting information is available online: Simulated particle events and comparison of models, Simulated cavity disturbance, Signal processing, Table of results and GNP specifications.